# Bayesian Multivariable Bidirectional Mendelian Randomization

### v1

@@ 4 Real Data Analysis
In this section, we demonstrate the effectiveness of the proposed method by applying it to two real-world genomic datasets: the skeletal muscle samples from the GTEx v7 dataset and the B-cell samples from the OneK1K dataset. For each dataset, we consider individuals who have complete information of gene expressions, SNPs, and relevant covariates (e.g., sex and age).
We apply our algorithm to infer the causal gene regulatory networks, with the associated uncertainty quantified by the posterior probabilities of edge inclusion, and assess the presence and the structure of latent confounders. Because horizontal pleiotropy may link any SNP to multiple genes in real tissues, we use the MR.RGM+ variant, allowing the model to select the relevant SNP–gene pairs. We run the proposed MCMC with 50,000 iterations, a burn-in of 10,000, and thinning every 10 iterations, yielding 4,000 posterior samples.
4.1 Application to GTEx v7 Skeletal Muscle Tissue Data
@@
The GTEx project is a comprehensive resource designed to study the relationship between genetic variation and gene expression across multiple human tissues. We focus on the skeletal muscle samples from GTEx v7, which consists of 332 individuals with both genotype and gene expression data available. Our analysis centers around the mechanistic target of rapamycin (mTOR) signaling pathway, a key regulator of cell growth and metabolism, which has been widely studied in both physiological and pathological contexts. We select 18 genes that are well-established components or regulators of the mTOR signaling cascade:
@@
We extract normalized expression levels of these 18 genes for the 332 individuals, resulting in a $332\times 18$ gene expression matrix. We utilize the signif_variant_gene_pairs file provided by GTEx and identify 62 SNPs that show significant association with at least one of the 18 genes. The resulting genotype (instrument) matrix has dimensions $332\times 62$. In addition to gene expression and SNPs, we control for two individual-level covariates: sex and age. The sex variable is coded as binary (male/female), while age is discretized into ordinal bins: 20–29 $\to$ 1, 30–39 $\to$ 2, …, 70–79 $\to$ 6.
@@ Results.
Figure  8 shows the estimated causal network. For simplicity, we only display causal relationships for which the posterior inclusion probabilities (PIPs) are over 0.85 and confounding relationships for which the PIPs are over 0.5. Blue arrows denote directed causal edges (double-headed where bidirectional), and orange curved links indicate latent confounding between gene pairs. Edge color shading indicates the PIP, reflecting the estimation uncertainty.
Figure 8: GTEx skeletal muscle mTOR signaling network. For clarity, we only display causal edges with inclusion probability $\geq 0.85$ and confounding edges with inclusion probability $\geq 0.50$. Blue arrows: causal edges (double-headed if bidirectional). Oranger curved edges: latent confounding links. Edge color shading reflects posterior support.
In Tables 1 and 2, we highlight a subset of biologically plausible causal relationships (gene regulations) and confounding structure, respectively. Generally, they align well with known molecular interactions within the mTOR signaling pathway and related regulatory cascades in muscle tissue. These results are obtained without imposing any prior knowledge about the underlying network topology, demonstrating the power of the proposed method in discovering interpretable gene regulatory relationships from genomic data.
@@
MTOR $\rightarrow$ S6K 0.871 One of the most prominent and canonical interactions in the mTORC1 pathway is the phosphorylation of ribosomal protein S6K by mTOR. This activation is essential for promoting protein synthesis and cell growth ma2009mTOR .
MTOR $\rightarrow$ PDK1 0.950 MTOR has been shown to regulate PDK1 activity, particularly under nutrient and growth factor stimulation. This regulatory axis is crucial for coordinating upstream Akt signaling and mTORC1 activation manning2007akt .
@@
VHL $\rightarrow$ INSULIN RECEPTOR 0.984 The tumor suppressor VHL regulates hypoxia-inducible factors (HIFs), which in turn affect insulin sensitivity and receptor expression, linking VHL to metabolic regulation Gunton2020HIF .
PHLPP1/2 $\rightarrow$ PKCA 0.987 PHLPP phosphatases dephosphorylate PKC isoforms, including PKCA, playing a role in signal termination downstream of PI3K/Akt Newton2010PKC .
Tel2 $\rightarrow$ TSC2 0.878 Tel2, part of the TTT complex, stabilizes PIKK proteins such as mTOR. By influencing mTOR stability, Tel2 indirectly affects TSC2 through mTOR-mediated regulation takai2007tel2 .
MLST8 $\rightarrow$ PRAS40 0.899 MLST8 is a core component of mTORC1 and mTORC2. Through its scaffolding function, it modulates assembly of complexes that regulate PRAS40, a known mTORC1 inhibitor VanderHaar2007PRAS40 .
@@
Table 1: A few key gene regulatory relationships identified from the GTEx v7 skeletal muscle dataset, along with their posterior inclusion probabilities (PIPs) and biological validation.
@@
MTORTSC2 0.525 MTOR and TSC2 are part of the same regulatory axis wherein TSC2 negatively regulates mTORC1 activity. Their interaction is modulated by AMPK and insulin signaling, making their co-expression sensitive to metabolic state and upstream energy-sensing signals, a plausible source of shared confounding saxton2017mtor .
PDK1S6K 0.501 In human skeletal muscle, both PDK1 and S6K are activated downstream of insulin/PI3KAktmTOR signaling. While PDK1 can directly phosphorylate S6K1, their correlation at the systems level is more likely driven by confounding through shared upstream inputs, particularly insulin- and growth factor–mediated PI3K activity. Thus, rather than reflecting a direct causal dependency, their association in muscle tissue may arise from coordinated regulation of the anabolic signaling network alessi1998pdk1 .
Folliculin/BHDPRAS40 0.503 Both Folliculin (BHD) and PRAS40 negatively regulate mTORC1 in response to metabolic cues and are phosphorylated in response to AMPK/Akt signaling. Their activities intersect through nutrient-sensing regulatory feedbacks, suggesting coordinated regulation and shared latent influences tsun2013folliculin .
MTORFKBP12 0.501 FKBP12 is a well-characterized binding partner of MTOR, crucial for rapamycin-mediated inhibition of mTORC1. Their expression is often correlated under rapamycin treatment and nutrient signaling, indicative of potential shared upstream regulatory programs sabatini1994raft1 .
Table 2: A few biologically plausible confounding structures inferred from the GTEx v7 skeletal muscle dataset, presented with their posterior inclusion probabilities (PIPs) and supporting biological validation.
4.2 Application to OneK1K B Cell Data
@@
For these genes, we extract normalized cell-type-level gene expressions across all 891 individuals, resulting in a $891\times 66$ gene expression matrix. SNP-gene marginal association scores are obtained from the OneK1K study. We retain 847 SNPs significantly associated with at least one of the 66 genes, producing an instrument matrix of dimension $891\times 847$. We also include as covariates sex (male/female) and age (discretized: <<30 as 1, 30–39 as 2, 40–49 as 3, …, 70–79 as 6, and 80+ as 7).
@@ Results.
In Tables 3 and 4, we highlight a subset of biologically plausible causal relationships (gene regulations) and confounding structure, respectively. Many of them align with well-established regulatory mechanisms in B cell development, signal transduction, and immune response modulation. They capture not only unidirectional regulatory relationships but also feedback loops that reflect the dynamic nature of BCR signaling.
@@
PYK2 $\rightarrow$ SYK 0.981 PYK2 (also known as PTK2B) is a calcium-sensitive non-receptor tyrosine kinase that plays a regulatory role in immune cell signaling. In B cells, PYK2 is activated downstream of integrin and BCR engagement and has been shown to facilitate the recruitment and activation of SYK by phosphorylating ITAMs and forming signaling complexes with SYK and other adaptors. This upstream role positions PYK2 as a modulator of SYK-mediated signaling cascades, particularly in integrin-enhanced or adhesion-dependent B cell responses Tse2009 .
CBL $\rightarrow$ DOK1 0.994 CBL, an E3 ubiquitin ligase, is recruited to signaling complexes downstream of the BCR where it associates with tyrosine-phosphorylated DOK1. This interaction promotes ubiquitination and degradation of DOK1, modulating its role as a negative regulator of Ras and PI3K signaling. Through this regulation, CBL fine-tunes BCR signaling by limiting DOK1’s inhibitory effects on MAPK and survival pathways, ensuring balance between activation and attenuation of B cell responses Lupher1999Cbl .
PIP3 $\rightarrow$ AKT 0.883 In B cells, PIP3 generated by PI3K recruits AKT to the membrane via its PH domain, where it becomes activated by phosphorylation. This is central to cell survival and metabolic regulation So2012 .
CD19 $\rightarrow$ PI3K 0.949 CD19 serves as a coreceptor for BCR signaling and amplifies signal transduction by recruiting PI3K. It directly binds the p85 regulatory subunit of PI3K upon phosphorylation, enhancing PIP3 production and facilitating activation of downstream effectors such as AKT and BTK in B cells Inabe2002 .
ERK1/2 $\rightarrow$ JUN 0.927 ERK1/2, part of the MAPK cascade, phosphorylates c-JUN, a component of the AP-1 transcription factor complex. In activated B cells, this phosphorylation increases JUN transcriptional activity, promoting expression of genes involved in proliferation, differentiation, and survival Roskoski2012 .
PI3K $\rightarrow$ AKT 0.857 The PI3KAKT signaling axis is central in B cell biology. Upon activation by coreceptors or cytokines, PI3K catalyzes production of PIP3, which recruits and activates AKT. This supports survival, proliferation, and metabolic adaptation Okkenhaug2003 .
@@
Table 3: A few key gene regulatory relationships identified from the OneK1K B-cell dataset, along with their posterior inclusion probabilities (PIPs) and biological validation.
@@
MEK1/2ERK1/2 0.498 In B cells, both MEK1/2 (MAP2K1/2) and ERK1/2 are co-activated by RasRaf signaling following BCR stimulation. Their correlation likely reflects confounding via shared RasRaf inputs that simultaneously activate MEK1/2 and ERK1/2 kolch2000 .
@@
MEKKJNK 0.494 MEKK (MAP3K1) and JNK (MAPK8/9) are key kinases in the MAPK signaling cascade in B cells. Their activation is strongly coordinated through shared upstream regulators, notably the CBM (CARMA1–BCL10–MALT1) complex and PKCB-mediated BCR signaling. The correlation between MEKK and JNK likely reflects confounding by common upstream effectors, as both respond to antigen receptor stimulation and stress signals Davis2000 .
@@
EZRINHS1 0.508 EZRIN links membrane proteins to the actin cytoskeleton, while HS1 regulates actin remodeling in hematopoietic cells. Both proteins are co-regulated during immune synapse formation in B cells through calcium and PI3K signaling. Their correlation likely arises from confounding by cytoskeletal remodeling pathways Harwood2011CytoskeletonBcell .
@@
PLCY2VAV 0.491 PLCY2 hydrolyzes PIP2 to trigger calcium release, while VAV is a GEF that activates Rac and cytoskeletal rearrangements. Both are simultaneously recruited to the BCR signalosome via SYK and BLNK, suggesting that their association reflects confounding through this shared scaffold Weber2008 .
@@
SHIPFGR2B 0.492 SHIP is an inositol phosphatase recruited by the inhibitory receptor FGR2B. Their correlation is expected since FGR2B engagement recruits SHIP via ITIM phosphorylation by LYN. Thus, their co-regulation is driven by shared LYN kinase activity, creating a confounding structure Barlev2022 .
@@
CREBMEF2C 0.487 CREB is activated downstream of PKA/ERK, while MEF2C responds to CAMK and MAPK signals. Both transcription factors are co-activated by BCR-induced calcium influx and MAPK cascades, producing correlated activity through shared upstream inputs rather than direct interaction Khiem2008 .
@@
RACPYK2 0.486 RAC is a small GTPase that controls actin dynamics, while PYK2 is a focal adhesion kinase activated by calcium and integrin signaling. In B cells, both are regulated downstream of VAV and calcium signals during immune synapse formation. Their association likely reflects confounding by BCR-driven adhesion and cytoskeletal pathways Tse2009 .
@@
Table 4: A few biologically plausible confounding structures inferred from the OneK1K B-cell dataset, presented with their posterior inclusion probabilities (PIPs) and supporting biological validation.

### v2

@@ 4 Real Data Analysis
In this section, we demonstrate the effectiveness of the proposed method by applying it to two real-world genomic datasets: the skeletal muscle samples from the GTEx v77 dataset and the B-cell samples from the OneK11K dataset. For each dataset, we discard individuals who do not have complete information of gene expressions, SNPs, or relevant covariates (e.g., sex and age).
We apply our algorithm to infer the causal gene regulatory networks, with the associated uncertainty quantified by the posterior probabilities of edge inclusion, and assess the presence and the structure of latent confounders. Because horizontal pleiotropy may link any SNP to multiple genes in real tissues, we use the MR.RGM+ variant, allowing the model to select the relevant SNP–gene pairs. We run the proposed MCMC with 50,00050{,}000 iterations, a burn-in of 10,00010{,}000, and thinning every $10$ iterations, yielding 4,0004{,}000 posterior samples.
Beyond edge-level posterior inclusion probabilities, we further summarize higher-order regulatory structure using the NetworkMotif function of our R package. This function computes posterior probabilities for user-specified network motifs—such as feedback loops, feedforward loops, and cascades—by quantifying the proportion of posterior samples in which all constituent edges are simultaneously present. This network motif-level summary provides a principled way to quantify the uncertainty of biologically meaningful regulatory patterns that extend beyond pairwise interactions.
4.1 Application to GTEx v77 Skeletal Muscle Tissue Data
@@
The GTEx project is a comprehensive resource designed to study the relationship between genetic variation and gene expression across multiple human tissues. We focus on the skeletal muscle samples from GTEx v77, which consists of $332$ individuals with both genotype and gene expression data available. Our analysis centers around the mechanistic target of rapamycin (mTOR) signaling pathway, a key regulator of cell growth and metabolism, which has been widely studied in both physiological and pathological contexts. We select $18$ genes that are well-established components or regulators of the mTOR signaling cascade:
@@
We extract normalized expression levels of these $18$ genes for the $332$ individuals, resulting in a $332\times 18$ gene expression matrix. We utilize the signif_variant_gene_pairs file provided by GTEx and identify $62$ SNPs that show significant association with at least one of the $18$ genes. The resulting genotype (instrument) matrix has dimensions $332\times 62$. In addition to gene expression and SNPs, we control for two individual-level covariates: sex and age. The sex variable is coded as binary (male/female), while age is discretized into ordinal bins: 20202929 $\to$ 11, 30303939 $\to$ 22, …, 70707979 $\to$ 66.
@@ Results.
Figure 10 shows the estimated causal network. For simplicity, we only display causal relationships for which the posterior inclusion probabilities (PIPs) are over $0.85$ and confounding relationships for which the PIPs are over $0.5$. Blue arrows denote directed causal edges (double-headed where bidirectional), and orange curved links indicate latent confounding of gene pairs. Within each edge type, lighter lines indicate lower PIP.
Figure 10: GTEx skeletal muscle mTOR signaling network. For clarity, we only display causal edges with inclusion probability $\geq 0.85$ and confounding edges with inclusion probability $\geq 0.50$. Blue arrows: causal edges (double-headed if bidirectional). Orange curved edges: latent confounding links. Edge color shading reflects posterior support.
In Tables 3 and 4, we highlight a subset of biologically plausible causal relationships (gene regulations) and confounding structure, respectively. Generally, they align well with known molecular interactions within the mTOR signaling pathway and related regulatory cascades in muscle tissue. These results are obtained without imposing any prior knowledge about the underlying network topology, demonstrating the power of the proposed method in discovering interpretable gene regulatory relationships from genomic data.
@@
MTOR $\rightarrow$ S6K $0.871$ One of the most prominent and canonical interactions in the mTORC1 pathway is the phosphorylation of ribosomal protein S6K by mTOR. This activation is essential for promoting protein synthesis and cell growth Ma and Blenis (2009).
MTOR $\rightarrow$ PDK1 $0.950$ MTOR has been shown to regulate PDK1 activity, particularly under nutrient and growth factor stimulation. This regulatory axis is crucial for coordinating upstream Akt signaling and mTORC1 activation Manning and Cantley (2007).
@@
VHL $\rightarrow$ INSULIN RECEPTOR $0.984$ The tumor suppressor VHL regulates hypoxia-inducible factors (HIFs), which in turn affect insulin sensitivity and receptor expression, linking VHL to metabolic regulation Gunton (2020).
PHLPP1/2 $\rightarrow$ PKCA $0.987$ PHLPP phosphatases dephosphorylate PKC isoforms, including PKCA, playing a role in signal termination downstream of PI3K/Akt Newton (2010).
Tel2 $\rightarrow$ TSC2 $0.878$ Tel2, part of the TTT complex, stabilizes PIKK proteins such as mTOR. By influencing mTOR stability, Tel2 indirectly affects TSC2 through mTOR-mediated regulation Takai et al. (2007).
MLST8 $\rightarrow$ PRAS40 $0.899$ MLST8 is a core component of mTORC1 and mTORC2. Through its scaffolding function, it modulates assembly of complexes that regulate PRAS40, a known mTORC1 inhibitor Vander Haar et al. (2007).
@@
Table 3: A few key gene regulatory relationships identified from the GTEx v77 skeletal muscle dataset, along with their posterior inclusion probabilities (PIPs) and biological validation.
@@
MTORTSC2 $0.525$ MTOR and TSC2 are part of the same regulatory axis wherein TSC2 negatively regulates mTORC1 activity. Their interaction is modulated by AMPK and insulin signaling, making their co-expression sensitive to metabolic state and upstream energy-sensing signals, a plausible source of shared confounding Saxton and Sabatini (2017).
PDK1S6K $0.501$ In human skeletal muscle, both PDK1 and S6K are activated downstream of insulin/PI3KAktmTOR signaling. While PDK1 can directly phosphorylate S6K1, their correlation at the systems level is more likely driven by confounding through shared upstream inputs, particularly insulin- and growth factor–mediated PI3K activity. Thus, rather than reflecting a direct causal dependency, their association in muscle tissue may arise from coordinated regulation of the anabolic signaling network Alessi et al. (1998).
Folliculin/BHDPRAS40 $0.503$ Both Folliculin (BHD) and PRAS40 negatively regulate mTORC1 in response to metabolic cues and are phosphorylated in response to AMPK/Akt signaling. Their activities intersect through nutrient-sensing regulatory feedbacks, suggesting coordinated regulation and shared latent influences Tsun et al. (2013).
MTORFKBP12 $0.501$ FKBP12 is a well-characterized binding partner of MTOR, crucial for rapamycin-mediated inhibition of mTORC1. Their expression is often correlated under rapamycin treatment and nutrient signaling, indicative of potential shared upstream regulatory programs Sabatini et al. (1994).
Table 4: A few biologically plausible confounding structures inferred from the GTEx v77 skeletal muscle dataset, presented with their posterior inclusion probabilities (PIPs) and supporting biological validation.
In addition to individual regulatory edges, we examined higher-order regulatory motifs within the inferred mTOR signaling network. We identify a feedback loop involving PDK1, VHL, and PKCA (Figure 12(a)). PDK1 acts as a central kinase upstream of multiple metabolic pathways, while VHL regulates hypoxia-inducible signaling that intersects with insulin and mTOR pathways (Gunton, 2020). PKCA participates in insulin-responsive signaling and modulates downstream mTOR activity (Newton, 2010). The simultaneous presence of these interactions suggests a regulatory feedback architecture integrating nutrient sensing, hypoxia response, and kinase signaling. The posterior probability of this motif is $0.854$.
We further identify a feedforward loop in which PDK1 regulates S6K both directly and indirectly through VHL (Figure 12(b)), consistent with coordinated control of protein synthesis downstream of insulin and growth factor signaling (Balendran et al., 1999; Ma and Blenis, 2009). This feedforward structure exhibits a posterior probability of $0.724$.
[…]
4.2 Application to OneK11K B Cell Data
@@
For these genes, we extract normalized cell-type-level gene expressions across all $891$ individuals, resulting in a $891\times 66$ gene expression matrix. SNP-gene marginal association scores are obtained from the OneK11K study. We retain $847$ SNPs significantly associated with at least one of the $66$ genes, producing an instrument matrix of dimension $891\times 847$. We also include as covariates sex (male/female) and age (discretized: <<$30$ as $1$, 30303939 as $2$, 40404949 as $3$, …, 70707979 as $6$, and $80+$ as $7$).
@@ Results.
In Tables 5 and 6, we highlight a subset of biologically plausible causal relationships (gene regulations) and confounding structure, respectively. Many of them align with well-established regulatory mechanisms in B cell development, signal transduction, and immune response modulation. They capture not only unidirectional regulatory relationships but also feedback loops that reflect the dynamic nature of BCR signaling.
@@
PYK2 $\rightarrow$ SYK $0.981$ PYK2 (also known as PTK2B) is a calcium-sensitive non-receptor tyrosine kinase that plays a regulatory role in immune cell signaling. In B cells, PYK2 is activated downstream of integrin and BCR engagement and has been shown to facilitate the recruitment and activation of SYK by phosphorylating ITAMs and forming signaling complexes with SYK and other adaptors. This upstream role positions PYK2 as a modulator of SYK-mediated signaling cascades, particularly in integrin-enhanced or adhesion-dependent B cell responses Tse et al. (2009).
CBL $\rightarrow$ DOK1 $0.994$ CBL, an E3 ubiquitin ligase, is recruited to signaling complexes downstream of the BCR where it associates with tyrosine-phosphorylated DOK1. This interaction promotes ubiquitination and degradation of DOK1, modulating its role as a negative regulator of Ras and PI3K signaling. Through this regulation, CBL fine-tunes BCR signaling by limiting DOK1’s inhibitory effects on MAPK and survival pathways, ensuring balance between activation and attenuation of B cell responses Lupher et al. (1999).
PIP3 $\rightarrow$ AKT $0.883$ In B cells, PIP3 generated by PI3K recruits AKT to the membrane via its PH domain, where it becomes activated by phosphorylation. This is central to cell survival and metabolic regulation So and Fruman (2012).
CD19 $\rightarrow$ PI3K $0.949$ CD19 serves as a coreceptor for BCR signaling and amplifies signal transduction by recruiting PI3K. It directly binds the p85 regulatory subunit of PI3K upon phosphorylation, enhancing PIP3 production and facilitating activation of downstream effectors such as AKT and BTK in B cells Inabe and Kurosaki (2002).
ERK1/2 $\rightarrow$ JUN $0.927$ ERK1/2, part of the MAPK cascade, phosphorylates c-JUN, a component of the AP-1 transcription factor complex. In activated B cells, this phosphorylation increases JUN transcriptional activity, promoting expression of genes involved in proliferation, differentiation, and survival Roskoski (2012).
PI3K $\rightarrow$ AKT $0.857$ The PI3KAKT signaling axis is central in B cell biology. Upon activation by coreceptors or cytokines, PI3K catalyzes production of PIP3, which recruits and activates AKT. This supports survival, proliferation, and metabolic adaptation Okkenhaug and Vanhaesebroeck (2003).
@@
Table 5: A few key gene regulatory relationships identified from the OneK11K B-cell dataset, along with their posterior inclusion probabilities (PIPs) and biological validation.
@@
MEK1/2ERK1/2 $0.498$ In B cells, both MEK1/2 (MAP2K1/2) and ERK1/2 are co-activated by RasRaf signaling following BCR stimulation. Their correlation likely reflects confounding via shared RasRaf inputs that simultaneously activate MEK1/2 and ERK1/2 Kolch (2000).
@@
MEKKJNK $0.494$ MEKK (MAP3K1) and JNK (MAPK8/9) are key kinases in the MAPK signaling cascade in B cells. Their activation is strongly coordinated through shared upstream regulators, notably the CBM (CARMA1–BCL10–MALT1) complex and PKCB-mediated BCR signaling. The correlation between MEKK and JNK likely reflects confounding by common upstream effectors, as both respond to antigen receptor stimulation and stress signals Davis (2000).
@@
EZRINHS1 $0.508$ EZRIN links membrane proteins to the actin cytoskeleton, while HS1 regulates actin remodeling in hematopoietic cells. Both proteins are co-regulated during immune synapse formation in B cells through calcium and PI3K signaling. Their correlation likely arises from confounding by cytoskeletal remodeling pathways Harwood and Batista (2011).
@@
PLCY2VAV $0.491$ PLCY2 hydrolyzes PIP2 to trigger calcium release, while VAV is a GEF that activates Rac and cytoskeletal rearrangements. Both are simultaneously recruited to the BCR signalosome via SYK and BLNK, suggesting that their association reflects confounding through this shared scaffold Weber et al. (2008).
@@
SHIPFGR2B $0.492$ SHIP is an inositol phosphatase recruited by the inhibitory receptor FGR2B. Their correlation is expected since FGR2B engagement recruits SHIP via ITIM phosphorylation by LYN. Thus, their co-regulation is driven by shared LYN kinase activity, creating a confounding structure Barlev et al. (2022).
@@
CREBMEF2C $0.487$ CREB is activated downstream of PKA/ERK, while MEF2C responds to CAMK and MAPK signals. Both transcription factors are co-activated by BCR-induced calcium influx and MAPK cascades, producing correlated activity through shared upstream inputs rather than direct interaction Khiem et al. (2008).
@@
RACPYK2 $0.486$ RAC is a small GTPase that controls actin dynamics, while PYK2 is a focal adhesion kinase activated by calcium and integrin signaling. In B cells, both are regulated downstream of VAV and calcium signals during immune synapse formation. Their association likely reflects confounding by BCR-driven adhesion and cytoskeletal pathways Tse et al. (2009).
@@
Table 6: A few biologically plausible confounding structures inferred from the OneK11K B-cell dataset, presented with their posterior inclusion probabilities (PIPs) and supporting biological validation.
Next, we characterize higher-order regulatory motifs in the B cell receptor (BCR) signaling network inferred from the OneK11K dataset. A prominent feedback loop involves PLCY2, RAC, and ERK1/2 (Figure 12(d)), reflecting the well-established coupling between calcium signaling, cytoskeletal reorganization, and MAPK activation during B cell activation (Oh-Hora and Rao, 2008; Roskoski, 2012). Such feedback interactions are known to stabilize signaling dynamics and support sustained transcriptional responses. This motif is strongly supported by the posterior, with a probability of $0.926$.
[…]
(a) GTEx feedback motif (PDK1VHLPKCA; $p=0.854$).
(b) GTEx feedforward motif (PDK1VHLS6K; $p=0.724$).
(c) GTEx cascade motif (VHLPKCAMTOR; $p=0.931$).
(d) OneK1K feedback motif (PLCY2RACERK1/2; $p=0.926$).
(e) OneK1K feedforward motif (BCRP70S6KIKK; $p=0.918$).
(f) OneK1K cascade motif (LYNRACERK1/2; $p=0.949$).
Figure 12: Motif-level summaries inferred using the NetworkMotif function. Each subfigure shows a feedback loop, feedforward loop, or cascade, whose uncertainty is quantified by its posterior probability calculated from posterior samples (the proportion of posterior draws in which all edges comprising the motif are simultaneously present).